BE12-8 In Eastwood Co., capital balances are Irey $40,000 and Pedigo $50,000. The part- ners share income equally. Vernon is admitted to the firm with a 45% interest by an invest- ment of cash of $58,000. Journalize the admission of Vernon.
Answer #1

Admission of a partner journal entries

Cash account. Dr 58000

To Vernon's Capital account . 58000

(Being capital brought in by Vernon in cash)

Vernon's Capital account. Dr. 17500

To Irey capital account. (17500*1/2) 8750

To Pedigo capital account .(17500*1/2) 8750

(Being goodwill brought in by Vernon adjust through the capital accounts)

Computation of goodwill:

Cash brought in by Vernon=$58000

45% share in company= (40000+50000)*45%= $40500

Goodwill=58000-40500= $17500

Note: the goodwill will be distributed among old partners equally because nothing had been mentioned in the question regarding the sacrifice made by old partners individually to accommodate Vernon. Therefore it is safe to assume that they sacrifice their shares equally to give 45% share to Vernon.